OpenStudy (anonymous):

find y' using implicit differentiation: cosx+y^1/2=4

OpenStudy (anonymous):

implicit differentiation wrt x: treat every y as an unknown function y(x). this means that you always use the chain rule: answer is: \[-\sin(x) + (1/2)y^{-\frac{ 1 }{ 2 }}*y' = 4\] to find y', isolate it. in this case it's rather simple: \[y' = \frac{ 4 + sinx }{ (1/2)y^{-\frac{ 1}{ 2 } }} = (8 + 2sinx)\sqrt{y}\]

OpenStudy (anonymous):

this is incorrect

OpenStudy (anonymous):

what is the correct answer?